# How Do Document Parsers Break? Auditing Structural Vulnerability in Document Intelligence

### 19 May 2026 v1

@@ Abstract
Document Layout Analysis (DLA) pipelines provide structured page representations for retrieval-augmented generation, long-document question answering, and other document intelligence systems, yet their robustness evaluation remains largely area-centric. We identify this Footprint Bias and propose a lightweight output-level auditing framework that decouples probe construction, policy-driven targeting, and structure-aware diagnosis. The framework combines Block-level Structural Loss Rate (B-SLR), granularity-aware exposure descriptors, and pathway attribution to analyze where perturbations interact with layout structure and how failures propagate. Across MinerU and PP-StructureV3 on 1,000 pages, affected area weakly tracks perturbation-induced OCR instability ($R^{2}{=}0.384/0.110$), whereas B-SLR aligns much more closely with it ($R^{2}{=}0.727/0.916$). Exposure descriptors further separate occlusion- and topology-dominant pathways, and small structurally targeted probes cause downstream QA/retrieval degradation comparable to larger-footprint perturbations. These results shift DLA robustness evaluation from footprint-based stress testing toward structure-aware vulnerability auditing.
@@
Yue Chen1,*, Yihao Wang1,*, Ziyi Tang1, Keze Wang1,† 1Sun Yat-sen University *Equal contribution.  Correspondence: kezewang@gmail.com
@@ 1 Introduction
Document Layout Analysis (DLA) converts visually organized pages into structured representations of text, tables, figures, and other layout elements. These representations serve as structural input for document intelligence systems such as retrieval-augmented generation over visually rich documents (Ueda et al., 2026), financial document reasoning (Zhao et al., 2024a), and medical record understanding (Cottet et al., 2026). Errors at this stage—including boundary corruption, spurious block merges, and misplaced outputs—can propagate as corrupted context, misaligned evidence, and unreliable downstream predictions.
@@
Despite progress on clean benchmarks and realistic evaluation settings, DLA robustness is still commonly assessed through aggregate degradation under corruption. Existing protocols often parameterize perturbation severity by global corruption magnitude or affected pixel footprint, while reporting damage through metrics such as CER or detection accuracy (Hendrycks and Dietterich, 2019; Michaelis et al., 2020; Chen et al., 2024; Du et al., 2025; Zhou et al., 2026). Such evaluations tell us whether performance declines, but reveal little about how or why parsing fails structurally. Moreover, annotation-tied diagnostics provide limited support for lightweight auditing on unlabeled document collections.
@@
Figure 1: Footprint Bias in DLA robustness evaluation: (a) a large-area perturbation may cause limited error, while (b) a small structural probe can trigger greater parsing failure.
A central limitation is what we term the Footprint Bias: the tendency to infer perturbation severity from pixel footprint alone. As illustrated in Figure 1, small-footprint structural probes can become disproportionately damaging when they intersect sensitive layout regions. Robustness therefore depends not only on how much of the page is perturbed, but also on where the perturbation falls and how the failure propagates through the parsing process.
This gap motivates a more diagnostic view of DLA robustness. Since naturally occurring parser failures are difficult to isolate and compare directly, we use controlled perturbations as counterfactual probes on real pages. Rather than asking only whether output quality deteriorates, we ask three finer-grained questions: (Q1) where structural identity is lost, (Q2) at what granularity perturbation exposure becomes predictive of failure, and (Q3) which pathway—direct physical occlusion or topology-level disruption—drives the degradation.
To address the resulting diagnostic problem, we propose a lightweight, output-level tripartite auditing framework that decouples perturbation generation from structural diagnosis. It defines a shared probe space and a policy space for attributable probe selection without modifying the audited parser, and a diagnostic space that combines observable terminal signals with structure-aware failure analysis. In particular, Block-level Structural Loss Rate (B-SLR) compares clean and perturbed parser outputs, enabling structural-loss auditing without manual element annotations. Experiments on two representative parsers show that structural diagnostics align with perturbation-induced OCR instability and downstream QA/retrieval failures more faithfully than footprint-based severity. Together, these results turn DLA robustness evaluation from footprint-based stress testing into structure-aware vulnerability auditing for document intelligence pipelines.
@@
  • We identify the Footprint Bias, showing that pixel footprint is an unreliable proxy for structural damage in DLA robustness evaluation.
  • We propose a lightweight tripartite auditing framework that separates probe construction, policy-driven targeting, and structure-aware diagnosis.
  • We instantiate the framework with B-SLR, exposure descriptors, and pathway attribution, showing that structural diagnostics better align with OCR instability and downstream QA/retrieval failures.
@@ 2 Related Work
[…]
Perturbation footprint and placement sensitivity.
Visual perturbations are often controlled by magnitude, affected area, or local patch budget. Cutout and Random Erasing use erased area as a key augmentation strength (DeVries and Taylor, 2017; Zhong et al., 2020), while document augmentation tools such as Augraphy expose visual and spatial perturbation parameters for document images (Groleau et al., 2023). Patch-based robustness studies further show that location and composition matter: adversarial patches can remain effective under small spatial budgets (Brown et al., 2017), PatchMap reveals placement hot-spots (Kimhi et al., 2025), and IMPACT jointly optimizes patch shape, location, number, and content (Yang et al., 2026). These works suggest that perturbation strength depends not only on footprint, but also on where and how perturbations interact with visual structure. Our work brings this insight to document parsing, where small perturbations can become disproportionately damaging when they interact with layout structure.
Visual sensitivity beyond aggregate corruption.
A broader robustness literature also shows that model sensitivity cannot be fully characterized by a single global corruption score. Studies have linked robustness to frequency-dependent learning, Fourier-basis sensitivity, texture bias, and learned feature geometry (Rahaman et al., 2019; Tsuzuku and Sato, 2019; Geirhos et al., 2019; Yin et al., 2019; Wang et al., 2020; Li et al., 2025). We do not use these mechanisms as direct explanations of document parsing failures. Rather, they motivate a diagnostic stance: robustness evaluation should identify which visual-structural factors drive failure in the target system, rather than reducing sensitivity to a single global corruption score.
@@ 3 Problem Setting
$E=\mathcal{M}(I)=\{x_{i}\}_{i=1}^{N},$ (1)
where each parsed element is represented as $x_{i}=(b_{i},c_{i},t_{i})$, with bounding box $b_{i}$, category label $c_{i}\in\mathcal{C}_{5}$, and recognized text $t_{i}$. The canonical label space $\mathcal{C}_{5}$ contains five layout types: text, title, table, figure, and equation.
Given a probe configuration $P$, detailed in Section 4, we apply a visual perturbation operator $\mathbb{P}_{P}$ and obtain
$I^{\prime}=\mathbb{P}_{P}(I),\qquad E_{\mathrm{adv}}=\mathcal{M}(I^{\prime}).$ (2)
The audit evaluates the divergence between clean and perturbed parsing behavior:
$\mathcal{A}(E,E_{\mathrm{adv}};P,\mathcal{Y})=(\Delta,\mathcal{R}),$ (3)
where $\mathcal{Y}$ denotes optional benchmark references, such as ground-truth layout or text annotations when available. The terminal component $\Delta$ records externally visible degradation signals, such as OCR CER or detection mAP drop, using $\mathcal{Y}$ when required by the metric. These signals ground the audit in benchmark settings, but are not prerequisites for defining the structural component $\mathcal{R}$, which diagnoses where structural identity is lost, at what granularity perturbation exposure becomes predictive, and through which pathway failures occur. In our audit, CER is computed against the clean parser output to measure perturbation-induced OCR instability rather than absolute OCR correctness; detection degradation is measured by the per-image mAP@0.5 drop between clean and perturbed parser outputs.
Conventional robustness evaluation mainly reports $\Delta$. By introducing $\mathcal{R}$, our formulation shifts the goal from measuring aggregate degradation to diagnosing the structural effects of perturbation on the parsing process. This perturb-and-compare formulation is lightweight: it requires no parser retraining and reruns the same parser only on controlled perturbations of existing pages.
4 Auditing Framework
4.1 Tripartite Auditing Framework
As illustrated in Figure 2, we instantiate the auditing formulation with a tripartite framework:
@@ 4.1 Tripartite Auditing Framework
where $\mathcal{S}$ is a shared probe space, $\Pi$ is a policy space for selecting probes, and $\mathcal{D}$ is a diagnostic space for auditing the resulting parsing changes. This decomposition separates what perturbations can be applied, how they are selected for a document, and how their effects are measured, enabling existing parsers to be audited by changing probes or policies without modifying the parser itself.
Following Eq. 3, the diagnostic space is
$\mathcal{D}=\Delta\times\mathcal{R},$ (5)
where $\Delta$ denotes terminal degradation signals and $\mathcal{R}$ denotes structure-aware diagnostic information. In Figure 2, the left side instantiates controlled perturbation through $\mathcal{S}$ and $\Pi$, while the right side instantiates structural diagnosis through $\mathcal{D}$.
4.2 Controlled Perturbation
The perturbation side is designed to make naturally entangled failures attributable rather than merely observable. We abstract failure-inducing visual interactions into a shared probe space $\mathcal{S}$ and let different policies select from it, so differences in observed damage can be attributed to selection logic rather than search-space changes.
@@ 4.2 Controlled Perturbation
where $\mathcal{H}$ denotes geometry, $\mathcal{V}$ visual appearance, $\mathcal{B}$ composition behavior, and $\mathcal{T}$ placement strategy. Detailed probe definitions and parameter ranges are provided in Appendix A.1.
Given the shared probe space, a policy $\pi\in\Pi$ maps available document context $c$ to a selected probe:
$\pi:\;c\mapsto\hat{P}\in\mathcal{S}.$ (7)
Different policies use different levels of context, ranging from random and rule-based selection to LLM- and VLM-guided targeting. Implementation details are deferred to Appendix A.3.
4.3 Diagnostic Space
The diagnostic space $\mathcal{D}$ separates judging from explaining. The terminal component $\Delta$ records externally visible degradation signals, while the structural component $\mathcal{R}$ diagnoses how degradation arises inside the parsing process:
$\Delta=\bigl(\Delta_{\mathrm{ocr}},\Delta_{\mathrm{det}}\bigr),\quad\mathcal{R}=\bigl(\mathcal{R}_{\mathrm{fail}},\mathcal{R}_{\mathrm{grad}},\mathcal{R}_{\mathrm{path}}\bigr).$ (8)
We instantiate $\Delta_{\mathrm{ocr}}$ with OCR CER and $\Delta_{\mathrm{det}}$ with $\Delta\mathrm{mAP}$, the per-image mAP@0.5 drop between clean and perturbed parser outputs, following the metric definitions in Appendix B.4. CER serves as the primary terminal signal, while $\Delta\mathrm{mAP}$ provides a supplementary detection-side check. The three structural components correspond to the questions introduced in Section 1: $\mathcal{R}_{\mathrm{fail}}$ measures whether structural identity is lost, $\mathcal{R}_{\mathrm{grad}}$ describes at what granularity perturbation exposure becomes predictive of failure, and $\mathcal{R}_{\mathrm{path}}$ attributes failures to distinct propagation pathways.
4.4 Structural Diagnosis
We instantiate these three components as follows.
4.4.1 Structural Failure Measurement
We instantiate $\mathcal{R}_{\mathrm{fail}}$ by checking whether each clean parsed element preserves a valid geometry-text correspondence after perturbation. For a clean element $x=(b,c,t)\in E$, we assign its candidate perturbed counterpart by best-IoU lookup:
$m(x)=\arg\max_{x^{\prime}\in E_{\mathrm{adv}}}\mathrm{IoU}(x,x^{\prime}),$ (9)
where $\mathrm{IoU}(x,x^{\prime})$ denotes box overlap.
@@ 4.4.1 Structural Failure Measurement
$\mathrm{TextSim}(x,x^{\prime})=\frac{\left|\mathrm{LCS}\!\bigl(t(x),t(x^{\prime})\bigr)\right|}{\max\!\bigl(|t(x)|,\;|t(x^{\prime})|\bigr)},$ (10)
where $\mathrm{LCS}$ denotes character-level longest common subsequence. We write $x\sim x^{\prime}$ when both geometry and text are preserved:
$x\sim x^{\prime}$= $\bigl(\mathrm{IoU}(x,x^{\prime})\geq\tau_{\mathrm{iou}}\bigr)$ (11)
$\wedge\bigl(\mathrm{TextSim}(x,x^{\prime})\geq\tau_{\mathrm{text}}\bigr),$
@@
$\mathcal{R}_{\mathrm{fail}}:=\mathrm{B\mbox{-}SLR}=\frac{\left|\{x\in E:\neg(x\sim m(x))\}\right|}{|E|}.$ (12)
@@
4.4.2 Granularity-Aware Damage Descriptors
We instantiate $\mathcal{R}_{\mathrm{grad}}$ with five exposure descriptors that summarize how a probe interacts with page structure before terminal quality is evaluated. These descriptors span two practical settings: output-derived descriptors for reference-free auditing, and annotation-dependent descriptors for benchmark-side diagnosis when layout annotations are available. Let $\Omega$ denote the page domain, $S(P)\subseteq\Omega$ the perturbation support induced by probe $P$, $\Omega_{L}$ the union of annotated layout regions, $\partial\Omega_{L}$ the corresponding boundary support, $L$ the set of annotated layout elements, and $E$ the clean parser output.
@@ 4.4.2 Granularity-Aware Damage Descriptors
$\alpha(P;\Gamma)$ $=\frac{|S(P)\cap\Gamma|}{|\Gamma|},$ (13)
$\iota(P;\mathcal{X})$ $=\frac{1}{|\mathcal{X}|}\sum_{x\in\mathcal{X}}\mathbf{1}\!\bigl[b(x)\cap S(P)\neq\emptyset\bigr].$
@@
$\mathcal{G}_{\mathrm{area}}(P)$ $=\bigl(\alpha(P;\Omega),\alpha(P;\Omega_{L}),\alpha(P;\partial\Omega_{L})\bigr),$ (14)
$\mathcal{G}_{\mathrm{struct}}(P)$ $=\bigl(\iota(P;L),\iota(P;E)\bigr).$
Together, these two families instantiate $\mathcal{R}_{\mathrm{grad}}$. Concretely, $\mathcal{G}_{\mathrm{area}}$ corresponds to total occlusion ratio (TOR), area coverage ratio (ACR), and boundary pixel overlap (BPO), while $\mathcal{G}_{\mathrm{struct}}$ corresponds to block overlap count ratio (BOC) and element interference ratio (EIR). TOR and EIR are output-derived, whereas ACR, BPO, and BOC are annotation-dependent diagnostics used when layout annotations are available. Exact constructions of $\Omega_{L}$ and $\partial\Omega_{L}$ are deferred to Appendix B.1. This separation allows the core exposure analysis to remain output-level through TOR and EIR, while ACR, BPO, and BOC provide additional benchmark-side probes when annotations are available.
4.4.3 Failure Pathway Attribution
@@
$\mathcal{U}=\left\{x\in E:\neg\bigl(x\sim m(x)\bigr)\right\}$ (15)
@@
$\begin{cases}\mathcal{U}_{\mathrm{miss}}=\{x\in\mathcal{U}:\rho(x,P)\geq\eta_{\mathrm{occ}}\},\\[3.0pt] \mathcal{U}_{\mathrm{topo}}=\{x\in\mathcal{U}:\rho(x,P)<\eta_{\mathrm{occ}}\}.\end{cases}$ (17)

### 26 May 2026 v2

@@ Abstract
Document Layout Analysis (DLA) pipelines provide structured page representations for retrieval-augmented generation, long-document question answering, and other document intelligence systems, yet their robustness evaluation remains largely area-centric. We identify this Footprint Bias and propose ProSA, a lightweight output-level auditing framework that decouples controlled probing, policy-driven targeting, and structure-aware diagnosis. ProSA combines Block-level Structural Loss Rate (B-SLR), granularity-aware exposure descriptors, and pathway attribution to analyze where structural identity is lost, at what exposure granularity failures emerge, and how failures propagate. Across MinerU and PP-StructureV3 on 1,000 pages, affected area weakly tracks perturbation-induced OCR instability ($R^{2}{=}0.384/0.110$), whereas B-SLR aligns much more closely with it ($R^{2}{=}0.727/0.916$). Exposure descriptors further separate occlusion- and topology-dominant pathways, while matched-footprint structural probes cause much larger downstream QA/retrieval drops than area-matched erasure. These results shift DLA robustness evaluation from footprint-based measurement toward structure-aware vulnerability auditing.
@@
Yue Chen1,*, Yihao Wang1,*, Ziyi Tang1, Yongsen Zhen2, Keze Wang1,† 1Sun Yat-sen University 2Nanyang Technological University {cheny2639,wangyh357,tangzy27}@mail2.sysu.edu.cn, z.yongsensmile@gmail.com,  kezewang@gmail.com *Equal contribution.  Correspondence: kezewang@gmail.com
@@ 1 Introduction
Document Layout Analysis (DLA) converts visually organized pages into structured representations of text, tables, figures, and other layout elements, supporting document intelligence systems such as retrieval-augmented generation over visually rich documents (Ueda et al., 2026), financial document reasoning (Zhao et al., 2024a), and medical record understanding (Cottet et al., 2026). Errors at this stage—including boundary corruption, spurious block merges, and misplaced outputs—can propagate as corrupted context, misaligned evidence, and unreliable downstream predictions.
@@
Despite progress on clean benchmarks and realistic evaluation settings, DLA robustness evaluation remains largely area-centric. Existing protocols often parameterize perturbation severity by global corruption magnitude or affected pixel footprint, and report aggregate degradation through metrics such as Character Error Rate (CER) or detection accuracy (Hendrycks and Dietterich, 2019; Michaelis et al., 2020; Chen et al., 2024; Du et al., 2025; Ouyang et al., 2025). Such evaluations tell us whether performance declines, but reveal little about how or why parsing fails structurally, especially when lightweight auditing is needed on unlabeled document collections.
@@
Figure 1: Footprint Bias: a large perturbation may preserve structure, while a small structural probe can drop a text line and cause greater parsing damage.
A central limitation is what we term the Footprint Bias: the tendency to infer perturbation severity from pixel footprint alone. As Figure 1 shows, a broad erased region may corrupt local text while leaving the main paragraph structure largely intact. By contrast, a thin perturbation with a much smaller footprint can intersect a sensitive line or layout boundary, causing a missing-line failure in which the parser skips a line and connects non-adjacent text fragments as if they were continuous. The smaller perturbation can therefore be more damaging: it removes document evidence and disrupts the local reading flow despite occupying far fewer pixels.
Such mismatches can propagate downstream: a dropped line may remove answer-bearing evidence or corrupt retrieval/QA context, while a larger visible perturbation may leave block identity and reading order intact. They also expose why area-centric evaluation can mis-rank structural vulnerability. The key challenge is therefore not simply to measure degradation, but to explain structural vulnerability through three questions: (Q1) where structural identity is lost, (Q2) at what granularity perturbation exposure becomes predictive of failure, and (Q3) which pathway—direct physical occlusion or topology-level disruption—drives the degradation.
To address this diagnostic problem, we propose ProSA (Probe-guided Structure-aware Auditing), a lightweight output-level framework for auditing structural vulnerability in DLA pipelines. ProSA instantiates the audit as a tripartite decomposition $\mathcal{F}=(\mathcal{S},\Pi,\mathcal{D})$, where $\mathcal{S}$ defines a shared space of controlled visual probes, $\Pi$ selects probes under different targeting policies, and $\mathcal{D}$ diagnoses the resulting clean–perturbed parsing divergence. By decoupling the probe space, selection logic, and diagnostic criteria, ProSA compares different targeting policies under a common action space without modifying the audited parser.
The diagnostic component is centered on B-SLR, an output-level signal that measures structural loss by comparing clean and perturbed parser outputs. Unlike footprint-based severity or terminal-only degradation scores, B-SLR checks whether each clean block preserves a valid geometry–text counterpart, capturing both missing geometry and lost textual identity under a unified structural criterion. Exposure descriptors and pathway attribution then explain at which exposure granularity these losses emerge and whether they arise from direct occlusion or topology-dominant disruption. Because the core diagnostics use the clean parse as structural reference, ProSA supports lightweight auditing on unlabeled documents while still incorporating benchmark annotations when available. Experiments show that these structural diagnostics better track OCR instability and downstream QA/retrieval drops than footprint-based severity.
@@
(1) We identify the Footprint Bias in DLA robustness evaluation, showing that pixel footprint is an unreliable proxy for structural damage and may mis-rank parser vulnerability.
(2) We propose ProSA, a lightweight output-level auditing framework that couples controlled probes, policy-guided selection, and structure-aware diagnosis.
(3) We instantiate ProSA with B-SLR, exposure descriptors, and pathway attribution, enabling annotation-light analysis that better tracks OCR and QA/retrieval drops.
@@ 2 Related Work
Robustness evaluation has moved from generic corruption suites and detection benchmarks (Hendrycks and Dietterich, 2019; Michaelis et al., 2020) to document-specific parsing settings such as RoDLA, DocPTBench, and OmniDocBench (Chen et al., 2024; Du et al., 2025; Ouyang et al., 2025). Recent document intelligence studies further show that OCR and layout quality can affect retrieval-augmented generation, financial reasoning, and medical document understanding (Ueda et al., 2026; Zhao et al., 2024a; Cottet et al., 2026; Zhang et al., 2025). However, these evaluations mainly report corruption-level, benchmark-level, or downstream degradation. They provide limited diagnosis of which parser-level structural changes, such as block loss, text-geometry mismatch, or topology disruption, actually drive the observed failures. Our work complements this line by auditing structural failure patterns directly in DLA outputs and relating them to downstream QA/retrieval behavior.
Perturbation footprint, placement, and black-box diagnosis.
Visual perturbations are often parameterized by magnitude, affected area, or patch budget. Cutout and Random Erasing use erased area as a key augmentation strength (DeVries and Taylor, 2017; Zhong et al., 2020), while document augmentation tools expose visual and spatial perturbation parameters for document images (Groleau et al., 2023). Patch-based robustness studies further show that location and composition matter: adversarial patches can remain effective under small spatial budgets (Brown et al., 2017), and recent analyses reveal placement hot-spots or jointly optimize patch shape, location, number, and content (Kimhi et al., 2025; Yang et al., 2026). Black-box explanation methods similarly perturb inputs to identify influential features or sensitive regions (Ribeiro et al., 2016; Lundberg and Lee, 2017; Petsiuk et al., 2018). Unlike these works, we audit structured DLA outputs rather than scalar predictions: each parsed element couples geometry, category, and text. This requires element matching, geometry-text preservation analysis, and occlusion/topology pathway diagnosis, while still requiring no access to parser internals.
@@ 3 Problem Setting
$E=\mathcal{M}(I)=\{e_{i}\}_{i=1}^{N},$ (1)
where each parsed element is represented as $e_{i}=(b_{i},c_{i},s_{i})$, with bounding box $b_{i}$, canonical category label $c_{i}\in\mathcal{C}_{5}=\{\text{text},\text{title},\text{table},\text{figure},\text{equation}\}$, and recognized text $s_{i}$.
Given a probe configuration $P$, we obtain a counterfactual page and its parser output:
$I^{\prime}=\operatorname{Perturb}(I;P),\qquad E_{\mathrm{adv}}=\mathcal{M}(I^{\prime}).$ (2)
We formulate auditing as comparing the clean and perturbed parsing outputs:
$(\Delta,\mathcal{R})=\operatorname{Audit}(E,E_{\mathrm{adv}};P,\mathcal{Y}),$ (3)
where $\mathcal{Y}$ denotes optional benchmark references when available. The terminal component $\Delta$ records externally visible degradation, including perturbation-induced OCR CER and detection-side $\Delta\mathrm{mAP}$, while the structural component $\mathcal{R}$ diagnoses where structural identity is lost, at what exposure granularity failure becomes predictive, and through which pathway it occurs. In our experiments, CER is computed against the clean parser output, and $\Delta\mathrm{mAP}$ is the per-image mAP@0.5 drop from clean to perturbed outputs against layout annotations.
This formulation extends conventional robustness evaluation beyond reporting $\Delta$: by introducing $\mathcal{R}$, the audit shifts from aggregate degradation measurement to structural perturbation diagnosis, without parser retraining or access to parser internals.
4 Probe-guided Structure-aware Auditing
As illustrated in Figure 2, we propose ProSA (Probe-guided Structure-aware Auditing), a lightweight output-level framework for auditing structural vulnerability in DLA pipelines. ProSA organizes the audit into two coupled components: a Policy-guided Prober, which constructs controlled perturbations, and a Structure-aware Auditor, which diagnoses the resulting clean–perturbed output divergence. Formally, we instantiate ProSA as
@@ 4 Probe-guided Structure-aware Auditing
where $\mathcal{S}$ is a shared probe space, $\Pi$ is a policy space for selecting probes, and $\mathcal{D}$ is a diagnostic space for auditing the resulting parsing changes. This decomposition separates what perturbations can be applied, how they are selected, and how their effects are measured, enabling DLA parsers to be audited without modifying the parser itself.
4.1 Component I: Policy-guided Prober
The Policy-guided Prober defines the controlled perturbation side of ProSA. It contains two coupled parts: a shared probe space $\mathcal{S}$ that defines admissible visual interventions, and a policy space $\Pi$ that selects probes under different targeting assumptions. Because all policies draw from the same $\mathcal{S}$, differences in observed damage can be attributed to selection logic rather than search-space changes.
Shared probe space.
@@ 4.1 Component I: Policy-guided Prober
where $\mathcal{H}$ denotes geometry, $\mathcal{V}$ visual appearance, $\mathcal{B}$ composition behavior, and $\mathcal{T}$ placement strategy. This factorization keeps the perturbation space modular: geometry controls the spatial support, appearance controls the visible pattern, behavior controls how the probe is composed with the page, and placement controls where it interacts with layout structure. Figure 5 in Appendix A.1 visualizes representative perturbation examples, while detailed probe definitions and parameter ranges are provided there.
Policy space.
Given the shared probe space, a policy $\pi\in\Pi$ maps available document context to a selected probe:
$\pi(\text{context})=\hat{P}\in\mathcal{S}.$ (6)
[…]
4.2 Component II: Structure-aware Auditor
Diagnostic space.
The Structure-aware Auditor instantiates the diagnostic space $\mathcal{D}$ by separating judging from explaining. The terminal component $\Delta$ records externally visible degradation signals, while the structural component $\mathcal{R}$ diagnoses how degradation arises inside the parsing process:
$\Delta$ $=\bigl(\Delta_{\mathrm{ocr}},\Delta_{\mathrm{det}}\bigr),$ (7)
$\mathcal{R}$ $=\bigl(\mathcal{R}_{\mathrm{fail}},\mathcal{R}_{\mathrm{exposure}},\mathcal{R}_{\mathrm{path}}\bigr).$
We instantiate $\Delta_{\mathrm{ocr}}$ with OCR CER and $\Delta_{\mathrm{det}}$ with $\Delta\mathrm{mAP}$, the per-image mAP@0.5 drop against layout annotations between the clean and perturbed outputs, following the metric definitions in Appendix B.4. CER serves as the primary terminal signal, while $\Delta\mathrm{mAP}$ provides a supplementary detection-side check. The three structural components correspond to the questions introduced in Section 1: $\mathcal{R}_{\mathrm{fail}}$ measures whether structural identity is lost, $\mathcal{R}_{\mathrm{exposure}}$ describes at what granularity perturbation exposure becomes predictive of failure, and $\mathcal{R}_{\mathrm{path}}$ attributes failures to distinct propagation pathways.
Structural failure measurement.
We instantiate $\mathcal{R}_{\mathrm{fail}}$ by checking whether each clean parsed element preserves a valid geometry-text correspondence after perturbation. For a clean element $e=(b,c,s)\in E$, we assign its candidate perturbed counterpart by best-IoU lookup:
$m(e)=\arg\max_{e^{\prime}\in E_{\mathrm{adv}}}\mathrm{IoU}(e,e^{\prime}),$ (8)
where $\mathrm{IoU}(e,e^{\prime})$ denotes box overlap.
@@ 4.2 Component II: Structure-aware Auditor
$\mathrm{TextSim}(e,e^{\prime})=\frac{\left|\mathrm{LCS}\!\bigl(s(e),s(e^{\prime})\bigr)\right|}{\max\!\bigl(|s(e)|,\;|s(e^{\prime})|\bigr)},$ (9)
where $\mathrm{LCS}$ denotes character-level longest common subsequence. We write $e\sim e^{\prime}$ when both geometry and text are preserved:
$e\sim e^{\prime}$= $\bigl(\mathrm{IoU}(e,e^{\prime})\geq\tau_{\mathrm{iou}}\bigr)$ (10)
$\wedge\bigl(\mathrm{TextSim}(e,e^{\prime})\geq\tau_{\mathrm{text}}\bigr),$
@@
$\mathcal{R}_{\mathrm{fail}}:=\mathrm{B\mbox{-}SLR}=\frac{\left|\{e\in E:\neg(e\sim m(e))\}\right|}{|E|}.$ (11)
@@
Granularity-aware exposure descriptors.
We instantiate $\mathcal{R}_{\mathrm{exposure}}$ with five exposure descriptors that summarize how a probe interacts with page structure before terminal quality is evaluated. These descriptors span two practical settings: output-derived descriptors for reference-free auditing, and annotation-dependent descriptors for benchmark-side diagnosis when layout annotations are available. Let $\Omega$ denote the page domain, $A(P)\subseteq\Omega$ the perturbation support induced by probe $P$, $\Omega_{L}$ the union of annotated layout regions, $\partial\Omega_{L}$ the corresponding boundary support, $L$ the set of annotated layout elements, and $E$ the clean parser output.
@@
$\operatorname{cover}(P,\Gamma)=\frac{|A(P)\cap\Gamma|}{|\Gamma|},$
$\operatorname{hit}(P,\mathcal{X})=\frac{1}{|\mathcal{X}|}\sum_{e\in\mathcal{X}}\mathbf{1}\!\bigl[b(e)\cap A(P)\neq\emptyset\bigr].$ (12)
@@
$\mathcal{G}_{\mathrm{area}}(P)=\bigl(\operatorname{cover}(P,\Omega),\operatorname{cover}(P,\Omega_{L}),$
$\qquad\operatorname{cover}(P,\partial\Omega_{L})\bigr),$
$\mathcal{G}_{\mathrm{struct}}(P)=\bigl(\operatorname{hit}(P,L),\operatorname{hit}(P,E)\bigr).$ (13)
Together, these two families instantiate $\mathcal{R}_{\mathrm{exposure}}$. Concretely, $\mathcal{G}_{\mathrm{area}}$ corresponds to total occlusion ratio (TOR), area coverage ratio (ACR), and boundary pixel overlap (BPO), while $\mathcal{G}_{\mathrm{struct}}$ corresponds to block overlap count ratio (BOC) and element interference ratio (EIR). TOR and EIR are output-derived, whereas ACR, BPO, and BOC are annotation-dependent diagnostics used when layout annotations are available. Exact constructions of $\Omega_{L}$ and $\partial\Omega_{L}$ are deferred to Appendix B.1. This separation allows the core exposure analysis to remain output-level through TOR and EIR, while ACR, BPO, and BOC provide additional benchmark-side probes when annotations are available.
Failure pathway attribution.
@@
$\mathcal{U}=\left\{e\in E:\neg\bigl(e\sim m(e)\bigr)\right\}$ (14)
@@
$\begin{cases}\mathcal{U}_{\mathrm{miss}}=\{e\in\mathcal{U}:\rho(e,P)\geq\eta_{\mathrm{occ}}\},\\[3.0pt] \mathcal{U}_{\mathrm{topo}}=\{e\in\mathcal{U}:\rho(e,P)<\eta_{\mathrm{occ}}\}.\end{cases}$ (16)